if a polygon is regular and one of the vertex angles is 165 degrees, how many sides must it have?
Answer by Alan3354(69443)

if a polygon is regular and one of the vertex angles is 165 degrees, how many sides must it have?
------------
Exterior angles = 180 - 165 = 15 degs
360/15 = 24 sides